Need answers

Chapter 11 Stockholders Equity: Paid-In Capital Stockholders equity Capital stock 7% cumulative preferred stock, $100 par value . . . . . . . . . . . . . . . . . . . . Common stock, $5 par value, 5,000,000 shares $ 15,000,000 20,000,000 44,000,000 64,450.6 Additional paid-in capital: Hetained earnings $143,450,000 From this information, compute answers to the following questions: a. How many shares of preferred stock have been issued? b. What is the total amount of the annual dividends to which preferred stockholders are entitled? c. What was the average issuance price per share of common stock? d. What is the amount of legal capital and the amount of total paid-in capital? e. What is the book value per share of common stock? f. Is it possible to determine the fair market value per share of common stock from the stock holders equity section above? Explain.
Answer #1

a.Number of shares of preferred stock that have been issued = Total Capital/Par value per Stock

= $15,000,000/$100

= 150,000 shares

b. The total amount of the annual dividends to which preferred stockholders are entitled = 7%*15,000,000

= $1,050,000

c. The average issuance price per share of common stock = (Par Value + Additional paid in Capital)/Number of shares issued

= (20,000,000+44,000,000)/4,000,000

= $16

d.Legal Capital = $15,000,000+20,000,000 = $35,000,000

Total paid in capital = Par value + Additional paid in capital

= $15,000,000 + 20,000,000+44,000,000

= $79,000,000

e. book value per share of common stock, assuming no dividends in arrears = (Par value + Additional paid in capital + retained earnings)/ Number of shares issued

= (20,000,000+ 44,000,000+ 64,450,000)/4,000,000

= $32.1125

f.No, it’s not possible to determine the fair market value per share of common stock from the stockholders' equity section above. Since the given information relates to book value, there’s no information about the market. Market value is very different from the book value, hence market value cannot be determined.